OpenStudy (anonymous):

If 1,520 Mega Watts of power, out of the 2,000 Mega Watts generated by a power utility, reach the destination, what percentage of the generated power reaches its destination?

OpenStudy (anonymous):

now deduct the 1,520 from 2000 and whatever u get divided by 2000X100

OpenStudy (anonymous):

ie 2000-1520=480/200x100=24

OpenStudy (anonymous):

200 or 2000x100

OpenStudy (anonymous):

i meann 2000

OpenStudy (anonymous):

did u get it

OpenStudy (anonymous):

lol. thanks cause I forgot how to do the simplest math. I think so I have a few more if you don't mind helping me really quick

OpenStudy (anonymous):

ok